OpenStudy (anonymous):

a) log4 (x+8) = log2 3 b) 3 ^2x= 5/3x c) log3 x = 2-log3 (x+8) d) log3 (2y-4) = 2 log3 2

OpenStudy (anonymous):

\[a) \log_{4} (x+8) = \log_{2} 3\]

OpenStudy (anonymous):

\[b) 3^{2x} = 5 / 3^{x}\]

OpenStudy (anonymous):

c) \[\log_{3} x=2-\log_{3} (x+8)\]

OpenStudy (anonymous):

\[\log_{3} (2y-4) = 2 \log_{3} 2 \]

OpenStudy (anonymous):

= log3 x2 + log3 5 = log3(5x2)

OpenStudy (anonymous):

@Callisto @FoolForMath @apoorvk @aroub

OpenStudy (apoorvk):